Space Stations

NASA International Space Station On-Orbit Status 14 October 2015

By Marc Boucher
Status Report
October 17, 2015
Filed under , ,
NASA International Space Station On-Orbit Status 14 October 2015
NASA International Space Station On-Orbit Status 14 October 2015.

In two weeks, NASA astronauts Scott Kelly and Kjell Lindgren will step outside the U.S. Quest airlock for the first of two maintenance spacewalks. The International Space Station is also being readied to host the next Orbital ATK Cygnus cargo mission set for early December.
Japanese astronaut Kimiya Yui has been servicing the two spacesuits Kelly and Lindgren will wear on the two six-hour spacewalks scheduled for Oct. 28 and Nov. 6. The spacewalkers will lubricate the tip of the International Space Station’s robotic arm, Canadarm2. He and Lindgren started their day, though, with eye checks for the ongoing Ocular Health study.

Kelly and Lindgren have also been preparing the Unity module where the Cygnus commercial cargo craft will be attached when it arrives in December after a 14-month hiatus. Kelly installed a Unity power adapter in the Destiny lab module then joined Lindgren to adjust power connectors inside Unity.

The three cosmonauts continued their routine maintenance tasks and science experiments in the station’s Russian segment. Flight Engineer Sergey Volkov explored crystal magnetism, while Oleg Kononenko and Mikhail Kornienko studied how a crew member adapts to motion during a spaceflight.

On-Orbit Status Report

Plant Gravity Sensing 2 (PGS2): Today Lindgren retrieved culture dishes from a Minus Eighty Degree Celsius Laboratory Freezer for ISS (MELFI) and treated the plants growing in them with a chemical reagent that induces a bioluminescence response that is dependent on the concentration of calcium ions in the plants. The Plant calcium concentrations have been shown to change in response to the direction of gravity: the plants emit photons when the culture dishes they are growing in are turned upside down so that the direction of gravity is changed relative to the plants. The reagent will be removed tomorrow with an absorbent and on Thursday the culture dishes will be turned upside down and returned to a Photon Counting Unit (PCU) so that photon measurements can be taken. The PGS2 investigation supports the study of cellular formation of the gravity sensors and the molecular mechanism for gravity sensing in plants grown in microgravity conditions.

Ocular Health: Lindgren and Yui initiated their Flight Day 90 Ocular Health testing by performing vision and tonometry tests and taking blood pressure measurements. The Ocular Health protocol calls for a systematic gathering of physiological data to characterize the risk of microgravity-induced visual impairment/intracranial pressure in ISS crewmembers. Researchers believe that the measurement of visual, vascular and central nervous system changes over the course of this experiment and during the subsequent post-flight recovery will assist in the development of countermeasures, clinical monitoring strategies, and clinical practice guidelines.

SOLAR: Measurements continue to be taken for European Space Agency’s (ESA’s) SOLAR investigation during the current sun visibility window which began last Friday, October 9th. The goal of the SOLAR instruments is to measure solar spectral irradiance and variability.

Habitability: Today Kelly documented his recent observations related to human factors and habitability for the Habitability investigation. Habitability assesses the relationship between crew members and their environment in order to better prepare for future long-duration spaceflights to destinations, such as near earth asteroids and Mars. Observations recorded during 6 month and 1 year missions can help spacecraft designers determine how much habitable volume is required, and whether a mission’s duration impacts how much space crew members need.

ISS Extravehicular Mobility Unit (EMU) Umbilical (IEU) Checkout: As a follow up to yesterday’s IEU replacement, Yui succesfully performed an IEU activation and checkout. On September 23, the crew found a damaged seal on the coolant loop outlet port in the connector that connects the IEU to the EMU. Both IEUs were due to be changed out early next year due to life time expiration. The team elected to replace both expiring IEUs to support upcoming EMU operations.

Node 1 Nadir Berthing Port Cabling: Last Friday, the crew was unable to mate two pre-routed power cables due to interference resulting from connector backshell clocking. Today, the crew succesfully re-clocked the 90° connectors and completed the installation of the wire harnesses to the Node 1 Forward Overhead and Lab Aft Overhead Bulkheads. These cables will provide power to Node 1 Nadir Berthing Port and Galley Rack. The crew also removed the airlock y-jumper that was installed as an interim measure last Friday.

Extravehicular Activity (EVA) Preparation: Kelly has packed International Docking Adapter (IDA) Cables into an EVA cable bag. The cables will be installed on Pressurized Mating Adapter (PMA)-3 during the ISS Upgrades EVA currently scheduled for October 28th. The ISS Program plans to establish two docking ports on the US On-orbit Segment compatible with the International Docking System Standard (IDSS). In the future, the docking adapters will be installed onto the existing PMA-2 and PMA3 interfaces. Each adapter will permanently convert a passive Androgynous Peripheral Attachment System (APAS) docking interface into a passive IDSS interface.

Special Purpose Dexterous Manipulator (SPDM) Troubleshooting: Today, Robotics Controllers performed data collection on the SPDM Arm 2 Shoulder Yaw Joint utilizing a diagnostics software patch. The troubleshooting was conducted as a result of an anomaly that occurred on the joint last Thursday, 08 October. The data will now be analyzed by ground teams.

Today’s Planned Activities
All activities were completed unless otherwise noted.

HMS Visual Testing Activity
EMU – Packing Cable Bag
Hose prep for sample collection from EDV-ZV (EDV 1198) Replacement of urine receptacle (МП) and filter-insert (Ф-В) in АСУ.
АСУ Activation after Replacement Life Support System (СОЖ)
PGS2 – Equipment Gathering
HMS – Vision Questionnaire
ТКГ 429 (SM Aft) Transfers and IMS Ops
Soyuz 717 Samsung tablet charging – end
Ocular Health (OH) Blood Pressure Operations
Ocular Health (OH) – Tonometry Test Configuration
Ocular Health Experiment – Tonometry – Crew Medical Officer (CMO)
Ocular Health Experiment – Tonometry – Subject
Ocular Health (OH) Blood Pressure Operations
MOTOCARD. Experiment Ops
Test activation of Vozdukh Atmosphere Purification System Emergency Vacuum Valves [АВК СОА]
Ocular Health Experiment – Tonometry – Crew Medical Officer (CMO)
Ocular Health Experiment – Tonometry – Subject
MOTOCARD. Operator Assistance with the Experiment
EMU – Long Dryout
Post-Tonometry Stowage
EMU – Long Dryout
MOTOCARD. Photography of the Experiment Ops
XF305 Camcorder Settings Adjustment
PGS2-ELT2 – Cable Reconfig
PGS2. Sample Retrieval from MELFI
HMS – Vision Test
PGS2. Sample Treatment Run 1
Power down hardware downstream of NOD2 UOP2 per Plug in Plan (PiP)
HMS – Vision Questionnaire
СКПФ1 & СКПФ2 Dust Filter Changeout. Reflect in IMS and clean MRM1 Gas-Liquid Heat Exchanger (ГЖТ)
Acoustic Dosimeter Setup for Crew-worn Measurement Sessions
Copy CLS 10 from SSC laptop
HMS Defibrillator Inspection
EMU – Long Dryout
Private Medical Conference
Node 1 Data Cable Routing/Installation
24-hour ECG Monitoring (termination)
On MCC Go Demating telemetry connectors from КЛ-108А
KULONOVSKIY KRISTALL. Copy and download data to hard drive for return
Removal of КЛ-108А Radio transmitter module.
Checkout of Compound Specific Analyzer-Combustion (CSA-CP)
24-hour Holter monitoring (terminate)
ТКГ 429 (SM Aft) Transfers and IMS Ops
Installation of КЛ-121/122Ц module
EMU – Long Dryout
CONTUR-2. Installation of master arm with adapter on panel 418
CONTUR-2. Experiment Session Ops
IMS Delta File Prep
De-installation and installation of TVS Antenna Feeder
СОЖ maintenance
Evening Work Prep
Start EMU Battery Charge
CONTUR-2. Experiment Session Ops
On MCC Go Mating Telemetry Connectors to TVS Antenna Feeder after installation
HABIT – Experiment Ops
Plug in Plan (PiP) – Node 2 – Hardware Power-up downstream UOP1
Crew conference with Astronaut Office
CONTUR-2. De-installation of master arm with adapter from panel 418
CONTENT. Experiment Ops
HAM radio session from Columbus
Preparation of reports for Roscosmos site
ECON-M. Observations and Photography
URAGAN Observations and Photography

Completed Task List Items

PCS R17 Hard Drive Pre-Gather
Safety Video Followup
61P USOS cargo unpack [In Work]
AMS Blanket Prep
Photo survey of PMM1O1 A&B

Ground Activities
All activities were completed unless otherwise noted.

Powerdowns in support of Node 1 Nadir Berthing Port cable connects
SPDM Troubleshooting

Three-Day Look Ahead:

Thursday, 10/15: Fundoscope, REBA H/W Checkout, OFV Measurements, Magvector
Friday, 10/16: Fundoscope, Ocular Health, X2R14, IMAX, JEMAL Shielding, EVA LLB Battery Charge
Saturday, 10/17: Weekly Cleaning, Crew Off Duty

QUICK ISS Status – Environmental Control Group:

Component – Status
Elektron – On
Vozdukh – Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) – Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) – On
Carbon Dioxide Removal Assembly (CDRA) Lab – Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 – Operate
Major Constituent Analyzer (MCA) Lab – Shutdown
Major Constituent Analyzer (MCA) Node 3 – Operate
Oxygen Generation Assembly (OGA) – Process
Urine Processing Assembly (UPA) – Standby
Trace Contaminant Control System (TCCS) Lab – Off
Trace Contaminant Control System (TCCS) Node 3 – Full Up

SpaceRef co-founder, entrepreneur, writer, podcaster, nature lover and deep thinker.